OpenStudy (anonymous):

In a bike race, a rider covers a 5-mi flat stretch of road at the speed of S mi/h. She then doubles her speed down a hill 1 mi long. Finally, she reduces her downhill speed by 12mi/h as she rides the last 3 mi of the race. What function gives the time T it takes the rider to finish the race in terms of S? I'm really confused. I don't want answers but I need someone to explain it

OpenStudy (anonymous):

\(r \cdot t=d\) solving for time we have \(t=\frac{d}{r}\)

OpenStudy (anonymous):

time = distance / rate

OpenStudy (anonymous):

can you help me work the problem?

OpenStudy (anonymous):

Start with the first part \(\displaystyle t=\frac{5}{S}+ \frac{1}{\frac{1}{2} S}+...\)

OpenStudy (anonymous):

It's sad that I know the formula, but I don't know how to work it out, isn't it? :/ Sorry.

OpenStudy (anonymous):

okay and hah me too but it's fine

OpenStudy (anonymous):

how do add the 1st part?

OpenStudy (anonymous):

5 divided by s + 1 divided by .5 * S.

OpenStudy (anonymous):

i'm so confused aha how do you do that?

OpenStudy (anonymous):

RATE * TIME = DISTANCE.

OpenStudy (anonymous):

i know that..i meant like the variable was confusing me

OpenStudy (anonymous):

Try this link, this guy explains how to solve this. Hope it works for you. http://www.connectionslivelesson.com/p85i34myxq3/?launcher=false&fcsContent=true&pbMode=normal